# A pedagogical approach to the black hole information issue

### v3

@@ II.2 Definition
The intuitive notion of a black hole is a region from which nothing can escape. To make this notion precise, one needs to describe what it means “to escape from” and to where it could be pertinent to “escape”. Notably, the idea that there is a restriction on the behavior of timelike and null curves is precisely captured by the concept of causally disconnected events. Namely, if there is no null or timelike curve connecting two events, then they are causally disconnected. Consequently, “to escape from” can be understood as the possibility of events to be causally connected. Additionally, the idea of where null curves could contemplate “escaping to” is given by future null infinity, $\mathscr{I}^{+}$. As discussed in Sec. II.1, in Minkowski spacetime all null curves end at $\mathscr{I}^{+}$, which means that, naturally, one expects that null curves would eventually reach $\mathscr{I}^{+}$, “escaping from” the rest of spacetime. In this sense, the definition of a black hole can be made precise in spacetimes possessing a region similar to $\mathscr{I}^{+}$. These are known as asymptotically flat spacetimes, as the “infinitely distant regions” have a Minkowski-like character. In other words, the “infinitely distant regions” to the past and the future in asymptotically flat spacetimes behave similarly to $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$.
In this manner, the black hole region of an asymptotically flat spacetime is defined as the set of events that are causally disconnected from $\mathscr{I}^{+}$. Physically, this can be interpreted as the fact that no electromagnetic wave emanating from regions inside the black hole can reach $\mathscr{I}^{+}$. In other words, nothing can escape from the black hole due to gravity, not even light. An example of this is given in fig. 6, which illustrates the conformal diagram of a spherically symmetric and electric uncharged black hole, known as a Schwarzschild black hole. The spacetime describing such a black hole is known as Schwarzschild spacetime, and the interpretation of its diagram is the following. Region II, highlighted in blue, is the black hole region. Region I is the exterior vacuum region, i.e., no classical energy distribution is present. Region III is the interior region of the spherically symmetric energy distribution that results in the black hole. Region II is delimited by the radial coordinate known as Schwarzschild radius,
@@
where $M_{\odot}$ is the Sun’s mass. The boundary of the black hole region, in this case characterized by the null curve obeying $r=r_{s}$, is known as the event horizon. Additionally, the thick white line at the top of the diagram represents the singularity inside the Schwarzschild black hole, a “pathological” region of the spacetime. The rest of the diagram boundary has the same interpretation as that of the conformal diagram of Minkowski spacetime.
Figure 6: Conformal diagram of Schwarzschild spacetime illustrating the exterior vacuum region (I), the black hole region (II) (highlighted in blue), and the collapsing energy distribution (III).
To visualize why region II is the black hole region, consider some examples of null and timelike curves, as shown in fig. 7.
@@
The timelike curve $\lambda_{2}$ is an example of the world-line of a massive particle that stays outside region II and that will end at $\mathcal{I}^{+}$. In contrast, the timelike curve $\lambda_{1}$ is an example of the world-line of a massive particle that enters region II, and its world-line ends not in $\mathcal{I}^{+}$, but rather, in the singularity. Similarly, the null curve $\gamma_{2}$ reaches $\mathscr{I}^{+}$, while $\gamma_{1}$ enters region II and eventually reaches the singularity. Indeed, any event in region II will not only not be able to communicate with events outside of it, but any massive particle or electromagnetic wave emanating from it will unavoidably reach the singularity. Notice that this is also the case for the energy distribution in region III. More precisely, after it collapses (i.e., reaches $r_{s}$), all the distribution will eventually reach the singularity.
In this manner, defining the black hole region as the set of events causally disconnected from $\mathscr{I}^{+}$ captures the idea of a region from which nothing can escape. However, the fact that $\mathscr{I}^{+}$ represents “infinitely distant regions” means that one would require complete knowledge of spacetime in order to analyze which events are causally connected to it. Due to this, it is convenient to restrict the definition of black holes to spacetimes that are globally hyperbolic at least outside the black hole and on the event horizon. With this, one would be able to study dynamical aspects of black holes in a covariant manner using Cauchy surfaces. Indeed, from figs. 8 and 9, it is easy to see that Schwarzschild spacetime is globally hyperbolic by using the same criterion as for Minkowski spacetime (see fig. 5). On the other hand, without this condition of global hyperbolicity, it is difficult to perceive a way to discuss a physically significant notion of time evolution of black holes.
@@
Figure 9: Sketch of the Finkelstein diagram corresponding to the conformal diagram shown in fig. 8. In this diagram, only null curves whose radial coordinate decrease over time are at $\pm 45\degree$, as shown by the “tilting” of the light cones as they approach $r=r_{s}$.
@@ II.3 Stationary black holes
Stationary black holes are those that do not change over time. Their analysis is pertinent mainly because of two reasons. First, one expects that all physical black holes (i.e., those resulting from gravitational collapse) [^6] In the following, we will only consider physical black holes. should be eventually described, to a good order of approximation, by a time independent configuration. Second, theorems in GR state that any stationary black hole can be described by only three parameters.
The first reason comes from the expectations regarding the product of gravitational collapse that results in a black hole. In particular, although details of the gravitational collapse can greatly affect the spacetime geometry, at sufficiently “late times” after its formation (as “measured” by Cauchy surfaces), a black hole is expected to reach a stationary configuration due to interaction with other sources of energy and the emission of gravitational waves (i.e., undulations of curvature propagating in spacetime). This conjecture is corroborated by the behavior of other systems, e.g., electromagnetic ones, in which a time dependent distribution of electric charges radiates away the higher order multipole moments and eventually “settles down” to a stationary distribution.
The second reason comes from the results known as the black hole uniqueness theorems. These theorems show that any stationary black hole must be described uniquely by its mass, $M$, angular momentum, $L$, and electric charge, $q$. The most general stationary black hole solution is known as a Kerr-Newman black hole described by the parameters $(M,L,q)$. Evidently, a Schwarzschild black hole is a nonrotating (i.e., $L=0$) and electric uncharged (i.e., $q=0$) Kerr-Newman black hole. Coupled with the conjecture that black holes should eventually reach a stationary configuration, the black hole uniqueness theorems leads to the conclusion that any black hole will eventually have only three degrees of freedom.
@@ III Semiclassical aspects of black holes
In this section, we will discuss the two main quantum effects that are pertinent to the conclusion of information loss in black holes. The first concerns the effective particle creation by black holes, known as the Hawking effect. The second is about an implication of entanglement between regions of spacetime in the framework of SG. Our starting point is to explain the concept of a particle content in QFT.
@@ III.1 Hawking effect
In QFT, a quantized field is the protagonist, being described by a collection of quantum harmonic oscillators in the spacetime. In this framework, what one would classically refer to as particles are now viewed as quanta or excitations of the oscillators describing the quantum field. As one would proceed in QM, the evaluation of a quantum-state particle content is done by the computation of the expectation value of the number operator. Consequently, dynamical processes that generate excitations of the quantum field can be understood as resulting in particle creation. This is the fundamental concept behind the Hawking effect.
In essence, an observer “near” $\mathscr{I}^{-}$ would identify an initial no particle state (i.e., one with no quanta or excitations) as the vacuum state, $|0\rangle_{\mathscr{I}^{-}}$. An example of such an observer is one at the early stages of its life, with a world-line given by $\lambda_{1}$ or $\lambda_{2}$ in fig. 7. Now, let $|0\rangle_{\mathscr{I}^{+}}$ denote the vacuum state for an observer “near” $\mathscr{I}^{+}$. An example of such an observer is one that remains outside the black hole at the late stages of its life, with a world-line also given by $\lambda_{2}$ in fig. 7. Notably, due to gravitational collapse and black hole formation, one finds that $|0\rangle_{\mathscr{I}^{-}}$ does not simply evolve to $|0\rangle_{\mathscr{I}^{+}}$, as illustrated in fig. 10.
@@
This means that observers at $\mathscr{I}^{+}$ perceive particles.
@@
where $k_{B}$ is Boltzmann’s constant and $\kappa$ is the black hole surface gravity. This result can be interpreted as stating that a Kerr-Newman black hole will effectively behave predominantly as a gray body [^7] The more detailed form of the spectrum contains gray body factors due to transmission probabilities., whose temperature is proportional to its surface gravity [^8] Although the local acceleration necessary to keep a massive particle still at the event horizon is divergent, the surface gravity is a finite quantity. This is because it is defined as the acceleration as measured from $\mathscr{I}^{+}$.. This effective particle creation by black holes is known as the Hawking effect, while the approximately thermal radiation predicted by it is referred to as Hawking radiation. As expected from the uniqueness theorems, a detailed evaluation of the spectrum for a Kerr-Newman black hole reveals that the emission of particles depends only on the parameters $(M,L,q)$. Also, the emission of particles with angular momentum and electric charge with the same sign as those of the black hole have a higher probability than those with opposite sign. Finally, the Hawking temperature of a Schwarzschild black hole is
$T\approx 6.18\;10^{-8}\left(\frac{M}{M_{\odot}}\right)^{-1}\;\text{K}.$ (4)
@@ III.2 Entanglement
Evidently, this is a pure composite state, because it is a product state of system $1$ (corresponding to particle 1) and system $2$ (corresponding to particle 2) that is represented by a single ket in the product Hilbert space. The fact that it is a pure state means that one has all the possible information about the composite system. However, by ignoring the degrees of freedom of system $1$ (i.e., by “tracing-out” its degrees of freedom in the state $|\psi\rangle$) one finds that system $2$ is described by the density operator
$\hat{\rho}=\frac{1}{2}|\uparrow_{2}\rangle\langle\uparrow_{2}|+\frac{1}{2}|\downarrow_{2}\rangle\langle\downarrow_{2}|,$ (6)
@@
i.e., a mixed state. This means that, by restricting one’s attention only to system $2$, there is now a probability distribution of states, and one cannot know for certain what state describes the individual system. It is in this sense that entanglement affects how much information one can acquire about the state of a single system which is a part of a larger one.
@@
In the context of SG, entanglement is an essential feature in light of what one would expect to be a physically acceptable class of states. In particular, since one is interested in describing backreaction effects (see fig. 2), it is of significance to understand properties that a state should posses in order to make sense of a notion of energy content. To understand this in more detail, consider the following line of reasoning.
@@
Recall that the ground state for a quantum harmonic oscillator has energy $\hbar\omega/2$. If one considers the infinite collection of quantum harmonic oscillators that describe a quantum field, then an associated vacuum state has a divergent energy content. Indeed, the nontrivial structure of the vacuum state in QFT leads to several effects that have been experimentally observed. However, no problems arise in the theoretical framework due to procedures that deal with the divergent quantities that appear (i.e., regularization and renormalization procedures). For example, in Minkowski spacetime, one performs a vacuum energy subtraction procedure using the natural vacuum state defined in it, the Minkowski vacuum, $|0\rangle_{M}$. This procedure allows one to give a finite notion of the energy content of a state, at least for states belonging to a certain class.
@@
Yet, such a procedure is not valid for general curved spacetime because there is no analog of the Minkowski vacuum. Nevertheless, one can select a class of states that satisfy the condition of having a divergent character similar to that of the Minkowski vacuum. States obeying this property are known as Hadamard states, and it can be argued that in order for a state to be physically acceptable it must be a Hadamard state. Indeed, it is possible to define a procedure that results in a finite energy content for Hadamard states Wald1994 ; Parker2009 .
@@
One of the main features of Hadamard states is that they imply entanglement over spacelike separated regions. In fig 11, we illustrate how this is of significance for black holes using two Cauchy surfaces, $\Sigma_{0}$ and $\Sigma_{1}$. Assuming that the initial state is a pure Hadamard state, then the field state at $\Sigma_{0}$ is also pure. The dynamical evolution of physical conditions from $\Sigma_{0}$ to any other Cauchy hypersurface will be unitary, and thus, will maintain its purity. For instance, for the “instant of time” $\Sigma_{1}$, the total field state is also pure. But, because $\Sigma_{1}$ intersects the black hole region, observers outside the black hole will have access only to a portion of $\Sigma_{1}$ (the dark red portion). Consequently, due to the entanglement implied by the Hadamard state, the field state inside the black hole “at a time” $\Sigma_{1}$ (the purple portion) will be entangled with the one outside. Hence, even though the field state at $\Sigma_{1}$ is pure, observers outside the black hole will not be able to acquire complete information about the state in the dark red portion. Obviously, they can always blame the black hole for concealing information about the quantum field, in the sense that information about the state inside the black hole could be accessed once one enters it. Nevertheless, information about the state inside the black hole will never be able to be shared with observers outside the black hole.
@@
Figure 11: Conformal diagram of a Schwarzschild spacetime illustrating two Cauchy hypersurfaces, $\Sigma_{0}$ (red) and $\Sigma_{1}$ (divided in two portions, dark red and purple).

### v4

@@ II.2 Definition
The intuitive notion of a black hole is a region from which nothing can escape. To make this notion precise, one needs to describe what it means “to escape from” and to where it could “escape”. Notably, the idea that there is a restriction on the behavior of timelike and null curves is precisely captured by the concept of causally disconnected events. Namely, if there is no null or timelike curve connecting two events, then they are causally disconnected. Consequently, “to escape from” can be understood as the possibility of events being causally connected. Additionally, the idea of where null curves could contemplate “escaping to” is given by future null infinity, $\mathscr{I}^{+}$. As discussed in sec. II.1, in Minkowski spacetime, all null curves end at $\mathscr{I}^{+}$, which means that, naturally, one expects that null curves would eventually reach $\mathscr{I}^{+}$, “escaping from” the rest of spacetime. In this sense, the definition of a black hole can be made precise in spacetimes possessing a region similar to $\mathscr{I}^{+}$. These are known as asymptotically flat spacetimes, as the “infinitely distant regions” have a Minkowski-like character. In other words, the asymptotic regions to the past and the future in asymptotically flat spacetimes behave similarly to $\mathscr{I}^{-}$ and $\mathscr{I}^{+}$.
In this manner, the black hole region of an asymptotically flat spacetime is defined as the set of events that are causally disconnected from $\mathscr{I}^{+}$. Physically, this can be interpreted as the fact that no information emanating from regions inside the black hole can reach $\mathscr{I}^{+}$. In other words, nothing can escape from the black hole due to gravity, not even light. An example of this is given in fig. 6, which illustrates the conformal diagram of a spherically symmetric and electric uncharged energy distribution that collapsed to form a black hole, known as a Schwarzschild black hole. The spacetime describing such a black hole is known as Schwarzschild spacetime, and the interpretation of its diagram is the following. Region II, highlighted in blue, is the black hole region. Region I is the exterior vacuum region, i.e., no classical energy distribution is present. Region III is the interior region of the spherically symmetric energy distribution that collapses to form a black hole. Region II is delimited by the radial coordinate defined by
$r=\left(\frac{A}{4\pi}\right)^{1/2},$ (2)
where $A$ is the black hole area. During the collapse, $A$ goes from zero to a maximum, such that the maximum value of the radial coordinate in eq. 2 is given by the Schwarzschild radius,
@@
where $M$ is the energy distribution total mass and $M_{\odot}$ is the Sun’s mass. The boundary of the black hole region is known as the event horizon. The thick white line at the top of the diagram represents the singularity inside the Schwarzschild black hole, a “pathological” region of spacetime. The rest of the diagram boundary has the same interpretation as that of the conformal diagram of Minkowski spacetime.
Figure 6: Conformal diagram of Schwarzschild spacetime illustrating the exterior vacuum region (I), the black hole region (II) (highlighted in blue), and the collapsing energy distribution (III). Note that the black hole region contains part of region III, and also an interior vacuum region.
To visualize why region II is the black hole region, consider some examples of null and timelike curves, as shown in fig. 7. The timelike curve $\lambda_{2}$ is an example of the worldline of a massive particle that stays outside region II, and that will end at $\mathcal{I}^{+}$. In contrast, the timelike curve $\lambda_{1}$ is an example of the worldline of a massive particle that enters region II, and its worldline ends not at $\mathcal{I}^{+}$, but rather, at the singularity. Similarly, the null curve $\gamma_{2}$ reaches $\mathscr{I}^{+}$, while $\gamma_{1}$ enters region II and eventually reaches the singularity. Indeed, any event in region II will not only be unable to communicate with events outside of it, but any massive particle, electromagnetic wave, or signal emanating from it will unavoidably reach the singularity. Note that this also holds for the energy distribution in region III. More precisely, after it collapses (i.e., reaches $r_{s}$), all the distribution will eventually reach the singularity.
@@
In this manner, defining the black hole region as the set of events causally disconnected from $\mathscr{I}^{+}$ captures the idea of a region from which nothing can escape. However, the fact that $\mathscr{I}^{+}$ represents an asymptotic region means that one would require complete knowledge of spacetime to analyze which events are causally connected to it. Due to this, it is convenient to restrict the definition of black holes to spacetimes that are globally hyperbolic at least outside the black hole and on the event horizon. With this, one can study the dynamical aspects of black holes in a covariant manner using Cauchy surfaces. Indeed, from figs. 8 and 9, it is easy to see that Schwarzschild spacetime is globally hyperbolic by using the same criterion as for Minkowski spacetime (see fig. 5). On the other hand, without this condition of global hyperbolicity, it is difficult to perceive a way to discuss a physically significant notion of time evolution of black holes.
@@
Figure 9: Sketch of the Finkelstein diagram corresponding to the conformal diagram shown in fig. 8. In this diagram, only null curves whose radial coordinate decreases over time are aligned with lines of slope $\pm 1$, as shown by the “tilting” of the light cones as they approach $r=r_{s}$.
@@ II.3 Stationary black holes
Stationary black holes are those that do not change over time. Their analysis is pertinent mainly for two reasons. First, one expects that all physical black holes (i.e., those resulting from gravitational collapse) [^8] In the following, we will only consider physical black holes. should be eventually described, to a good order of approximation, by a time independent configuration. Second, theorems in GR state that any stationary black hole can be described by only three parameters.
The first reason stems from expectations about the product of gravitational collapse which results in a black hole. In particular, although details of the gravitational collapse can greatly affect the spacetime geometry, at sufficiently “late times” after its formation (as “measured” by Cauchy surfaces), a black hole is expected to reach a stationary configuration due to interaction with other sources of energy and the emission of gravitational waves (i.e., undulations of curvature propagating in spacetime). This conjecture is corroborated by the behavior of other systems, e.g., electromagnetic ones, in which a time dependent distribution of electric charges radiates away the higher-order multipole moments and eventually “settles down” to a stationary distribution.
The second reason comes from the results known as the black hole uniqueness theorems. These theorems show that any stationary black hole must be described uniquely by its mass, $M$, angular momentum, $L$, and electric charge, $q$. The most general stationary black hole solution is known as a Kerr-Newman black hole, described by the parameters $(M,L,q)$. Evidently, a Schwarzschild black hole is a nonrotating (i.e., $L=0$) and electric uncharged (i.e., $q=0$) Kerr-Newman black hole. Coupled with the conjecture that black holes should eventually reach a stationary configuration, the black hole uniqueness theorems lead to the conclusion that any black hole will eventually have only three degrees of freedom.
@@ III Semiclassical aspects of black holes
In this section, we will discuss the two main quantum effects that are pertinent to the conclusion of information loss in black holes. The first concerns the effective particle creation by black holes, known as the Hawking effect. The second is about an implication of entanglement between regions of spacetime in the framework of SG. Our starting point is to explain the concept of particle content in QFT.
@@ III.1 Hawking effect
In QFT, a quantized field is the protagonist, being described by a collection of quantum harmonic oscillators in spacetime. In this framework, what one would classically refer to as particles are now viewed as quanta or excitations of the oscillators describing the quantum field. As one would proceed in QM, the evaluation of a quantum state particle content is done by the computation of the expectation value of the number operator. Consequently, dynamical processes that generate excitations of the quantum field can be understood as resulting in particle creation. This is the fundamental concept behind the Hawking effect.
Essentially, an observer “near” $\mathscr{I}^{-}$ would identify an initial no particle state (i.e., one with no quanta or excitations) as the vacuum state, $|0\rangle_{\mathscr{I}^{-}}$. An example of such an observer is one at the early stages of its life, with a worldline given by $\lambda_{1}$ or $\lambda_{2}$ in fig. 7. Now, let $|0\rangle_{\mathscr{I}^{+}}$ denote the vacuum state for an observer “near” $\mathscr{I}^{+}$. An example of such an observer is one that remains outside the black hole at the late stages of its life, with a worldline also given by $\lambda_{2}$ in fig. 7. Notably, due to gravitational collapse and black hole formation, one finds that $|0\rangle_{\mathscr{I}^{-}}$ does not simply evolve to $|0\rangle_{\mathscr{I}^{+}}$, as illustrated in fig. 10. This means that particles are created as witnessed by observers in the asymptotic future.
@@
where $k_{B}$ is Boltzmann’s constant and $\kappa$ is the black hole surface gravity. For a Schwarzschild black hole, the surface gravity reads
$\kappa=\frac{c^{4}}{4GM}\approx 10^{13}\left(\frac{M}{M_{\odot}}\right)^{-1}\;\text{m/s${}^{2}$}.$ (5)
This result can be interpreted as stating that a Kerr-Newman black hole will, effectively, behave predominantly as a gray body [^9] The more detailed form of the spectrum contains gray body factors due to transmission probabilities., whose temperature is proportional to its surface gravity [^10] Although the local acceleration necessary to keep a massive particle still at the event horizon is divergent, the surface gravity is a finite quantity. This is because it is defined as the acceleration as measured from $\mathscr{I}^{+}$.. This effective particle creation by black holes is known as the Hawking effect, while the approximate thermal radiation predicted by it is referred to as Hawking radiation. As expected from the uniqueness theorems, a detailed evaluation of the spectrum for a Kerr-Newman black hole reveals that the emission of particles depends only on the parameters $(M,L,q)$. Also, the emission of particles with angular momentum and electric charge with the same sign as those of the black hole has a higher probability than those with opposite sign.
@@ III.2 Entanglement
Evidently, this is a pure composite state, because it is represented by a single ket in the total Hilbert space. Additionally, it is an entangled state, as it cannot be written as a simple product state in the individual Hilbert spaces. Indeed, by ignoring the degrees of freedom of system $1$ (i.e., by “tracing-out” its degrees of freedom in the state $|\psi\rangle$) one finds that system $2$ is in a mixed state described by the density operator
$\hat{\rho}=\frac{1}{2}|\uparrow_{2}\rangle\langle\uparrow_{2}|+\frac{1}{2}|\downarrow_{2}\rangle\langle\downarrow_{2}|.$ (7)
@@
Notably, mixed states cannot be written as a ket in the Hilbert space. This means that, by restricting one’s attention only to system $2$, there is now a probability distribution of states, and one cannot know for certain what state describes the individual system. In this sense, entanglement affects how much information one can acquire about the state of a single subsystem.
@@
In the context of SG, entanglement is an essential feature in light of what one would expect to be a physically acceptable class of states. In particular, since one is ultimately interested in describing backreaction effects (see fig. 2), it is of significance to understand properties that a state should possess to make sense of a notion of energy content. To understand this in more detail, consider the following line of reasoning.
@@
Recall that the ground state for a quantum harmonic oscillator has energy $\hbar\omega/2$. If one considers the infinite collection of quantum harmonic oscillators that describe a quantum field, then an associated vacuum state has a divergent energy content. Indeed, the nontrivial structure of the vacuum state in QFT leads to several effects that have been experimentally observed. However, no problems arise in the theoretical framework due to procedures that address divergent quantities (i.e., regularization and renormalization). For example, in Minkowski spacetime, one performs a vacuum energy subtraction procedure using the natural vacuum state defined in it, the Minkowski vacuum, $|0\rangle_{M}$. This procedure allows one to give a finite notion of the energy content of a state, at least for states belonging to a certain class.
@@
Yet, such a procedure is not valid for a general curved spacetime because there is no analog of the Minkowski vacuum. Nevertheless, one can select a class of states that satisfy the condition of having a divergent character similar to that of the Minkowski vacuum. States obeying this property are known as Hadamard states, and it can be argued that for a state to be physically acceptable, it must be a Hadamard state. Indeed, it is possible to define a procedure that results in a finite energy content for Hadamard states Wald1994 ; Parker2009 .
@@
One of the main features of Hadamard states is that they imply entanglement over spacelike separated regions. In fig. 11, we illustrate how this is of significance for black holes using two Cauchy surfaces, $\Sigma_{0}$ and $\Sigma_{1}$. Assuming that the initial state is a pure Hadamard state, then the field state at $\Sigma_{0}$ is also pure. The dynamical evolution of physical conditions from $\Sigma_{0}$ to any other Cauchy hypersurface will be unitary, and thus, will maintain its purity. For instance, for the “instant of time” $\Sigma_{1}$, the total field state is also pure. But, because $\Sigma_{1}$ intersects the black hole region, observers outside the black hole will have access only to a portion of $\Sigma_{1}$ (the dark red portion). Consequently, due to the entanglement implied by the Hadamard state, the field state inside the black hole “at a time” $\Sigma_{1}$ (the purple portion) will be entangled with the one outside. Analogously to the spin case, the state outside the black hole will be mixed after one traces out degrees of freedom inside the event horizon. Evidently, they can always blame the black hole for concealing information about the quantum field, in the sense that information about the state inside the black hole could only be accessed once one enters it. Nevertheless, information about the state inside the black hole will never be able to be shared with observers outside the black hole. In other words, even though the spacetime is globally hyperbolic, observers outside the black hole will experience a mixed state simply because they do not have access to the black hole interior.
@@
Figure 11: Conformal diagram of Schwarzschild spacetime illustrating two Cauchy hypersurfaces, $\Sigma_{0}$ (red) and $\Sigma_{1}$ (divided in two portions, dark red (outside the black hole) and purple (inside the black hole)).